mathslover (mathslover):

Finding general term for : \[ \int \cfrac{px+q}{ax^2 + bx + c} dx \]

mathslover (mathslover):

@iambatman :-)

OpenStudy (anonymous):

My notifications are off so I just realized you tagged lol

mathslover (mathslover):

:-) Make the notifications on for me lol So, let us start this problem.

OpenStudy (anonymous):

Bleh it's a nasty one, but I guess you want to rewrite the integrand.

mathslover (mathslover):

I personally will like the coefficient of x in numerator and x^2 in denominator to become one. That is : \[\int \cfrac{ x + \cfrac{q}{p} }{x^2 + \cfrac{b}{a} x + \cfrac{c}{a} } \times \cfrac{a}{p} dx \]

mathslover (mathslover):

@ganeshie8 will also be able to help us! :-)

mathslover (mathslover):

Sorry, will it be p/a ? (just messed up everything :P )

OpenStudy (anonymous):

\[\int\limits \frac{ p(2ax+b) }{ 2a(ax^2+bx+c) }-\frac{ bp-2aq }{ 2a(ax^2+bx+c) }dx\] Expressed as a sum of fractions, you get that...

hartnn (hartnn):

\(\\ \text{9. To integrate }\quad \huge \frac{ax+b}{\sqrt{px^2+qx+r}} or \frac{ax+b}{px^2+qx+r} \\ \large express \quad ax+b=A\frac{d}{dx}(px^2+qx+r)+B \\ \text{then separate D.} \\~ \\~ \\~ \\~\) D = Denominator

mathslover (mathslover):

How did you get that? @iambatman I'm not sure that how did you do that.

mathslover (mathslover):

Oh.. hartnn got something from his tutorial for us.

OpenStudy (anonymous):

I think I may have done it wrong, just ignore it lol, just use what hartnn said., it's much easier.

mathslover (mathslover):

Yeah, what hartnn suggested was there in my book, but it was not well explained. I really don't know how they get these substitution ideas.

mathslover (mathslover):

I think I got how hartnn got that.

hartnn (hartnn):

you have to know the formulas. we have this formula : \(\\ \huge \text{In general,}\int \frac{f’(x)}{f(x)}\:dx=\ln|f(x)|+c \) which means if denominator is quadratic, we need to express numerator in terms of derivative of denominator to be able to use this formula

hartnn (hartnn):

also we have formulas for 1/(x^2+a^2) , 1/(a^2-x^2) and so on so even if we have a constant in numerator and quadratic in denominator, we should be able to solve it

OpenStudy (anonymous):

@mathslover what does it show on how they achieved the formula?

OpenStudy (anonymous):

I know of it, but I never knew how to derive it.

hartnn (hartnn):

we always try to express the given problem in terms of standard formulas. to be able to do that, we need to know all standard formulas which we can use. to know those, we need to practice good amount of problems on every known.

hartnn (hartnn):

on every known formula*

mathslover (mathslover):

Hartnn, yes, that was a great explanation even.

mathslover (mathslover):

I am getting it now... !

OpenStudy (anonymous):

Got it, thank you both :)!

mathslover (mathslover):

:-) Thank you @iambatman :)